OpenStudy (anonymous):

find the equation in slope-intercept form passing through the points (3,-12),(-4,2)

OpenStudy (mimi_x3):

Use the gradient formula to find m, that is m=-2 y = mx + b -12 = -2(3)+b -12 = -6 + b b = -6 y = -2x-6